a chemist needs to make 30 ounce of a 25% alcohol sulution by mixing together a 15% alcohal with 40% solution. how much of each solution should she use
Answer by checkley77(12844)

.40x+.15(30-x)=.25*30
.40x+4.5-.15x=7.5
.25x=7.5-4.5
.25x=3
x=3/.25
x=12 ounces of 40% solution is used.
30-12=18 ounces of 15% solution is used.
Proof:
.40*12+.15*18=.25*30
4.8+2.7=7.5
7.5=7.5